OpenStudy (samsan9):

solve the equation 6x^2+11x-35

OpenStudy (zzr0ck3r):

this is not an equation

OpenStudy (zzr0ck3r):

do you mean \(6x^2+11x-35=0\)?

OpenStudy (samsan9):

well just help me solve it atleast

OpenStudy (zzr0ck3r):

you cant solve unless you have it set equal to something

OpenStudy (samsan9):

i forgot to write =0

OpenStudy (zzr0ck3r):

\(\frac{-b\pm\sqrt{b^2-4ac}}{2a}\) for you \(a=6\\b=11\\v=-35\)

OpenStudy (anonymous):

Alternatively you can use factoring methods... for 6x2+11x−35=0

OpenStudy (samsan9):



OpenStudy (anonymous):

6x^2 + 21x - 10x -35 =0 and so on..

OpenStudy (samsan9):

where did you get -35 from?

OpenStudy (samsan9):

so i still don't get it

OpenStudy (anonymous):

is your question 6x2+11x−35=0

OpenStudy (samsan9):

yes

OpenStudy (anonymous):

6x2+11x−35=0 6x^2 + 21x - 10x -35 =0 3x(2x+7) -5(2x+7) =0 (3x-5)(2x+7) =0

OpenStudy (anonymous):

i hope you can solve for x now...

OpenStudy (samsan9):

yes i can thank you

OpenStudy (anonymous):

yw